# Emerging Quadrature Lattices of Kerr Combs

### v1

@@ 2 Results
where $c_{\mu}\equiv a_{\mu},d_{\mu}\equiv a_{-\mu}$ for $\mu>0$, and $d_{-1}\equiv c_{1},c_{-1}\equiv d_{1}$. We note that the higher order terms $\mathcal{O}\left(\alpha^{n}\right)$ add Bragg coupling and pair generation terms to $n$-nearest neighbours thereby not breaking the lattice structure. In this “folded” representation it is possible to show that if the detuning $\Delta\tilde{\omega}_{\mu}$ is symmetric around $\mu=0$, the bosonic quadratic lattice in Eq. LABEL:eq2 becomes a degenerate pseudo APT symmetric lattice which results in a eigenspectrum with eigenvalues that are purely real or purely imaginary (Methods 4.1).
To describe realistic squeezing behavior, the bosonic quadrature lattice of Eq. LABEL:eq2 must be finite, which naturally occurs due to group velocity dispersion (GVD) reducing the Bragg coupling coefficient for modes away from the center. If we define $\Delta\tilde{\omega}_{\mu}=\Delta\tilde{\omega}_{0}+\tilde{D}^{\prime}_{1}\mu+\tilde{D}_{2}\mu^{2}/2+\mathcal{O}(\mu^{3})$, then $\Delta\tilde{\omega}_{\mu}$ is symmetric if all the odd terms are zero. Note that ${D}^{\prime}_{1}(\tilde{D}^{\prime}_{1})$ is the (unitless) frequency difference between the cavity FSR and the Kerr comb’s frequency spacing. To illustrate the implication of different $D^{\prime}_{1}$ values, we calculate the quadrature eigenspectrum of the sub-threshold light generated by a 2-FSR Kerr comb with fixed $A_{\mu}$ profile and varying $D^{\prime}_{1}$. For concreteness, we use $\mathcal{R}$ and $D_{2}$ taken from our experimental system (see Methods 4.2). Figure 2(a) shows the symmetric case in which $D^{\prime}_{1}=0$, while Figs. 2b(c) are non-symmetric with $D^{\prime}_{1}/2\pi=3(8)MHz$. We adjust the overall comb power in the plots in Fig. 2 to be close to threshold for detuning $\Delta{\omega}_{0}=2g\sum_{\mu}\left|A_{\mu}\right|^{2}$. To compute the quadrature eigenspectrum we write Eq. LABEL:eq2 in the quadrature basis $(p_{1}...p_{N},q_{1}...q_{N})$ where $p_{i}$ and $q_{i}$ are the quadratures of mode $i$, and $N$ is the number of participating cavity modes. Since the transformation to quadratures is not a unitary operation, we obtain a non-Hermitian Hamiltonian $\mathcal{M}$. To study threshold behavior we introduce $\Gamma=\kappa/2I_{2N\times 2N}$, where $\kappa$ is the uniform intensity loss rate.
In Fig. 2 we analyze the eigenspectrum of $\mathcal{M}-\Gamma$ while varying the parameter $D^{\prime}_{1}$. We observe three distinct regimes. For $D^{\prime}_{1}=0$ (Fig. 2(a)), the eigenvalues are either purely real or have a real part that equals $-\kappa/2$ resulting from their initial purely real eigenvalue shifted uniformly by $\kappa/2$. Interestingly, due to the multimode nature of the system, even when the symmetry is removed (i.e. when $D^{\prime}_{1}\neq 0$) some states close to threshold remain with imaginary values of 0 (i.e the partially symmetric case, see Fig. 2(b)). This gradual decrease in the number of purely real eigenvalues or robustness is analyzed in Methods 4.2. For the highest $D^{\prime}_{1}$ in Fig. 2(c), the modes close to threshold all have a double peak structure (i.e., imaginary part is non zero).
Figure 2: Dispersion symmetry and the properties of supermode squeezed states. a.-c. Eigenvalues of the multimode states of a 2-FSR comb with different values of $D^{\prime}_{1}$: 0, 3 and 8 MHz respectively. For larger $D^{\prime}_{1}$ the purely real eigenvalues gradually become complex. Highlighted quadrature supermodes are those close to threshold. d.-g The frequency($\omega$)-dependent squeezing spectrum of the supermodes close to threshold (highlighted). Each line represents the supermode spectrum for a linearly swept intra cavity power, and the supermodes can be seen to narrow as they approach threshold. The supermode in f. does not narrow asymptotically because e. reaches threshold first. h.,i. The supermode composition as a function of cavity modes of the multimode squeezed states, for the supermodes in e. and f. respectively.
Each eigenvalue in Fig. 2(a-c) represents a multimode squeezed state, residing on the odd modes. As depicted in green in Fig. 1(a) the squeezing spectrum of each supermode uniformly evolves around each cavity mode [33]. The spectral structure around each cavity mode is frequency dependent squeezing, a concept which recently started to be utilized in metrology application through noise spectrum engineering [41].
To compute the multimode frequency-dependent squeezing structure, we calculate the supermodes of $\mathcal{M}-\Gamma$, using the Bloch-Messiah decomposition (details in Methods 4.2). The structure of the supermodes with the highest real part (gain) in Fig. 2(a-c) is presented in Fig. 2(d-g). Supermodes with the highest gain will dominate in total photon number over other supermodes and have the narrowest peaks, being closest to the optical parametric oscillator (OPO) threshold. The OPO threshold occurs when $Re(\lambda)=0$, where $\lambda$ denotes the eigenvalue of the supermode with the highest real part. The frequency spacing in the double peak structure is determined by the imaginary value of the corresponding supermode. Therefore, supermodes with imaginary value zero, exhibit a single peak structure (“un-detuned” from the frame of reference of the Kerr comb), while supermodes with imaginary value that is not zero exhibit double-peak spectrum (“detuned” from the frame of reference of the Kerr comb).
The partially symmetric case exhibits both double-peak and single peak evolution for different supermodes with high gain simultaneously, with frequency spacing between the peaks which is generally smaller than in the completely non-symmetric case. The different plotted lines in each spectral shape in Fig 2(d-g) represent the squeezing for different distances from threshold (by changing the overall power of the Kerr comb). Together with an intra-modal frequency-dependent structure, each supermode has a spectral structure in the cavity mode basis, shown for the partially symmetric case in Fig. 2(h-i). The bandwidth of these supermodes is predominantly determined by the values of $D^{\prime}_{1}$ and $D_{2}$.
Next, we aim to experimentally observe the frequency-dependent quadrature variance directly for the different cases described above using integrated microring resonators [42]. In this work, we utilize 4H-silicon carbide-on-insulator [43], which offers low losses and high third-order nonlinearity, enabling low-power operation. We excite a localized 2-FSR Kerr comb by injecting CW light near an avoided mode crossing, generating strong local anomalous dispersion. The comb is naturally phase-locked via cascaded parametric processes and centered within the telecommunications c-band (1545 nm). The characterization of the device parameters, including dispersion, is presented in Methods 4.3 and Methods 4.5).
We begin by characterizing the modal structure of the generated 2-FSR comb in its native basis. Using a custom-built single-photon optical spectrum analyzer (SPOSA)[7], which provides $>100$ dB of dynamic range with single-photon sensitivity, we simultaneously capture the photon populations above- and below-threshold, as shown in Fig. 3(a). Mapping out the two-photon correlation matrix of the below-threshold state can reveal the inter-modal connectivity in the Hamiltonian, and provide initial indications of the formation of the supermodes introduced in Fig. 1(b,c). We perform pairwise second-order photon correlation measurements, $g^{(2)}_{ij}(\tau)=\langle\hat{a}_{i}^{\dagger}(0)\hat{a}_{j}^{\dagger}(\tau)\hat{a}_{j}(\tau)\hat{a}_{i}(0)\rangle/{\langle\hat{a}_{i}^{\dagger}(0)\hat{a}_{i}(0)\rangle\langle\hat{a}_{j}^{\dagger}(\tau)\hat{a}_{j}(\tau)\rangle}$ on each pair of below-threshold modes in the comb, and the resulting matrix is shown in Fig. 3(b). The signature of supermodes is seen as pair-wise correlations between modes close in their distance from the center pump mode (consistent with the folded picture in Fig. 1(b)).
We proceed to demonstrate the transition between the non-symmetric regime and the symmetric regime through continuous tuning of the resonator dispersion properties. Gradually heating the microring changes the mode splitting, and therefore the threshold point, allowing us to vary $D^{\prime}_{1}$ (see Methods 4.3). Tuning our below-threshold comb above threshold, we observe through two-tone homodyne detection that adjusting the voltage of the heater changes the secondary comb frequency structure above threshold discontinuously (Fig. 3(c) and Methods 4.4). Figure 3(d-g) shows the above-threshold peaks for different applied voltages. We observe three distinct voltage regimes: In the first, the threshold presents a double peak structure with the peak separation of approximately 300 MHz (Fig. 3(d)), significantly larger than the cavity linewidth of  100 MHz, thus corresponding to the strongly non-symmetric regime of Fig. 2(c,g). Further tuning the voltage, we transition discontinuously to a regime in which the detuning is small, around 30 MHz, which corresponds to the partially symmetric regime of Fig. 2(b,e,f) followed by a transition to a single peak (corresponding to the symmetric case threshold crossing of Fig. 2(a,d)). The discontinuous change of the frequency spectrum with the voltage occurs because each regime in Fig. 2 corresponds to a different supermode crossing threshold, which is not a continuous process. Afterward, the system abruptly transitions to the partially-symmetric state again (Fig. 3(g)). This matches completely to the regimes described in Fig. 2.
Figure 3: 2-FSR Kerr comb and observation of the secondary threshold at different peak separation. a. Single-photon optical spectrum analyzer (SPOSA) measurement of photon populations in above threshold (red) and below threshold (green) modes of the 2-FSR state. b. The two-photon correlation ($g^{(2)}$) matrix of the below threshold light in the odd modes, showing the correlations generated by Bragg scattering and pair-generation in the quadrature lattice. c. Illustration of the homodyne measurement of the quadrature lattice, the LO is generated by an electro-optic modulator (EOM) and does not have the same spacing as the Kerr comb. d.-g. Homodyne measurements for four different applied heater voltages, showing the different regimes from symmetric to non-symmetric. The double peak is also observed with a single tone LO and with a LO located at other modes as well.
The above-threshold light shown in Fig. 3(d-g) presents only a partial picture of the eigenspectrum evolution, since it shows the single mode that reaches threshold first, without the below-threshold quadrature variance structure. To observe the competing multimode processes in the quantum regime, we also measure the below-threshold quadrature vacuum fluctuations (anti-squeezing), whose evolution precedes the formation of the above-threshold light shown in Fig. 3(d-g). To observe the below-threshold light, we reduce shot noise using on-chip and off-chip filters (Methods 4.3). Furthermore, we tune the power ratio of our two-tone LO to a ratio of $LO_{1}=10\cdot LO_{-1}$, which we found to maximize the overlap with the amplified quadrature supermode (Fig. 4(a)). Figure 4(b) shows our 2-FSR Kerr comb at the soliton crystal stage, with symmetry corresponding to the partially symmetric case (Fig. 2(b,e,f)). Below threshold, we observe both the single peak and double peak modes of Fig. 2(b,e,f) simultaneously. By further red-tuning the input pump frequency, we observe that while the detuned supermode is amplified and narrows further until it crosses threshold, the non-detuned supermode does not (Fig. 4(b)).
Focusing on the non-detuned state (Fig. 4(d,e)), we observe the oscillations (induced by the phase sweep) dip to shot noise. Squeezing is not observed due to out-coupling losses and incomplete overlap of the double tone LO with the supermode, but anti-squeezing is visible. The phase-dependent quadrature variance is consistent with the picture presented in Fig. 2(b,e,f), of the partially symmetric state, visible only through below-threshold quadrature measurements. In principle, the fine structure of the anti-squeezing is equivalent to that of the squeezing but more robust to losses making it useful for exploration. Using these states as a squeezed quantum resource would require overcoupling to the resonator and utilizing a LO with higher bandwidth, along with very low out-coupling losses.
Tuning more toward the symmetric state, we are able to measure the sub-threshold spectrum of the non-detuned supermode. We perform the same measurement as in Fig. 4(a) but for every two adjacent odd cavity modes simultaneously using a separate laser (Fig. 4(f)). The separate laser is not phased-locked to the pump; therefore, it does not show the quadrature phase dependency but does provides the quadrature mean variance value and spectrum. With this approach, we can resolve 3 separate peaks in each measurement (Fig. 4(g)). Two broad peaks represent the quadrature variance signals of two odd cavity modes, with comb leakage in between, which is narrow (since it is classical above-threshold light). Thus, retrieving each pair of neighbouring odd modes allows us to measure the general non detuned quadrature supermode structure, as shown in Fig. 4(h). The structure is similar in width to the observed correlations in Fig. 3(b). Similar analysis was performed on the detuned supermode giving similar multi-mode bandwidth since it is dictated by $D_{2}$ (Methods 4.6).
Figure 4: Below-threshold quadrature measurements: partially symmetric regime a. Scheme of the below-threshold homodyne measurement. Unlike the measurement in Fig. 3(d-g) here the ratio of LO tones is optimized and shot noise is filtered. b. Below-threshold quadrature oscillations induced by the phase shifter, showing the co-existence of two quadrature supermodes - non-detuned and detuned simultaneously, showing the partially symmetric regime directly. c. Tuning closer to resonance for the state in (b), we observe that the detuned state is increasing in gain while the non-detuned state remains weak, proving that there are two separate supermodes competing. d.-e. Zooming in on the frequency axis on the quadrature anti-squeezing oscillations of the un-detuned below threshold state. f.-g. Scheme and a measurement of mapping the non-detuned quadrature supermode by measuring 3 neighbouring modes with an additional laser. Each measurement produces two below-threshold quadrature variance together with a weak filtered comb line h. Quadrature variance of the non-detuned peaks, showing the profile of the non-detuned variance of the quadrature supermode.
The co-existense of quadrature supermodes and the robustness of single-peaked supermodes have significant implications for the process of comb formation and evolution: The two-peaked sub-threshold gain profile dictates the secondary comb frequency structure. It is revealing to understand how the different types of anti-squeezing structures affect the threshold from primary 2-FSR comb to a 1-FSR secondary comb. Before presenting the experimental results, we wish to gain further insights from the Lugiato–Lefever equations (LLE) [44] and simulate the classical non-linear evolution at the threshold (see LLE simulation details in Methods 4.5). By tuning the quality factors and dispersion in the simulation to impose a small $D^{\prime}_{1}$, we can simulate a non-detuned threshold to a secondary comb (Fig. 5(a-b)). In this type of threshold, the 2-FSR comb turns into a 1-FSR comb but does not form an rf beatnote as is generally the case [45]. Therefore, only after the first threshold to an un-detuned 1-FSR comb, a second threshold of a detuned comb occurs between the 1-FSR comb and an additional 1-FSR comb, which in-turn leads to an rf beatnote and transmission fluctuations Fig. 5(b).
Figure 5: Primary to secondary threshold behavior of the quadrature supermodes for different regimes. a-b. LLE simulation of a non detuned supermode that cross threshold and becomes a secondary 1-FSR comb. The simulation shows that it is followed by a second threshold in which a detuned state crosses threshold, which is then followed by the generation of light that causes an rf beatnote and transmission oscillations. c-d. The experimentally observed evolution of a non-detuned supermode into a secondary comb in the symmetric regime; c. shows the first threshold of a, d. is the second threshold that follows c and is detuned. e-f. A direct detuned threshold corresponding to the completely non-symmetric case. Unlike a-d there is only a single threshold before transmission oscillations. Plots e and f show the same threshold at different modes (+1 and +3).
Returning to our experiment, we follow the sub-threshold light with the same homodyne measurement as in Fig. 4(a). We take multiple snapshots of the quadrature structure during the process of threshold crossing. In Fig. 5(c), we observe how the non detuned sub-threshold anti-squeezed light crosses the threshold. The first threshold crossing results in a 1-FSR comb, and no rf beatnote. This is followed by a second quadrature supermode, this time detuned, crossing the threshold and giving rise to sub-combs (Fig. 5(d)). One can clearly see that after the threshold crossing of the un-detuned mode, the detuned mode still exists and keeps growing until it also crosses the threshold. Thus, our measurements reveal the mechanism for creating a 1-FSR comb from a 2-FSR primary comb without formation of an rf beatnote. The case of the un-detuned threshold contrasts with the case in which the detuned quadrature supermode reaches threshold first. In that case, as we show in two different measurements of the same quadrature supermode in Fig. 5(e,f), a comb with an rf beatnote can appear immediately after the first threshold. The two measurements are single tone homodyne measurements at different modes ($\mu=1,\mu=3$). Additionally, we present five more measurements in Methods 4.6 to show the uniformity of the quadrature variance structure along different cavity modes. This demonstrates how both types of quadrature supermodes result in different types of threshold behavior - one threshold vs two consecutive thresholds.
@@ 3 Discussion
In this work we demonstrated non-Hermitian lattice effects in photonic multimode quantum fluctuations described as photonic quadrature lattices. We used Kerr combs that generate photonic quadrature lattices to experimentally explore their non-Hermitian properties for the first time. By controlling the resonator dispersion via an integrated microheater, we were able to observe the three theoretically predicted regimes corresponding to different non-Hermitian lattice behavior. Namely, we observed the symmetric case, partially symmetric, and non-symmetric. Kerr combs allow scalability and rich complexity on a compact on-chip platform. By measuring the anti-squeezing of 2-mode projection of the quadrature supermode, we revealed the existence of detuned and robust non-detuned squeezed supermodes and explained how these form in the context of lattice symmetries. Exploring such fundamental properties paves the way for designing more complex lattice phenomena by injecting or probing different above-threshold states. Furthermore, understanding comb formation and the quantum noise of Kerr combs with our approach can lead to improved comb technology, particularly for combs that are governed by quantum noise which dictates the ultimate limit on comb stability. The concepts studied here form a bridge that will enable the exploitation of the rich non-Hermitian lattice physics in quadrature bosonic lattices to engineer complex quantum resources on-chip. The complex transport of quadratures offered by the Kerr effect and demonstrated in this work will lead to lattice phenomena such as non-Hermitian skin effect, Anderson localization, edge-states and topological phases. These phenomena would manifest as quadrature modes offering new opportunities to control quantum light.

### v3

@@ 2 Results
where $c_{\mu}\equiv a_{\mu},d_{\mu}\equiv a_{-\mu}$ for $\mu>0$, and $d_{-1}\equiv c_{1},c_{-1}\equiv d_{1}$. We note that the higher-order terms $\mathcal{O}\left(\alpha^{n}\right)$ introduce Bragg coupling and pair generation terms to the $n$-nearest neighbors, thereby preserving the lattice structure. In this ”folded” representation, if the detuning $\Delta\tilde{\omega}_{\mu}$ is symmetric around $\mu=0$, the Hamiltonian in Eq. LABEL:eq2 forms a degenerate pseudo APT-symmetric lattice. This results in an eigenspectrum with eigenvalues that are either purely real or purely imaginary (See Methods 4.5 for the symmetry analysis). As was discussed earlier, the real part of the eigenvalues corresponds to parametric amplification, while the imaginary part corresponds to oscillations. Consequently, when APT symmetry exists, extended supermodes which have gain close to threshold may not oscillate at all (purely real eigenvalues). However, in our microcombs, APT symmetry does not exist in general, mainly due to a mismatch between the repetition rate of the comb and the group-velocity $\Delta\tilde{\omega}_{\mu}$. Building on this realization, by adjusting the repetition rate of our microcomb, we can tune the oscillations of the supermode to absolute zero. We confirm the existence of extended non-oscillating supermodes generated by Kerr combs numerically (Methods 4.6), and experimentally as we will now show.
Tuning the repetition rate can be achieved by adjusting the micro-heater voltage $V$. This adjustment detunes the TE and TM modes differently, thereby altering the location of the mode crossing. Since the phase matching that initiates the 2-FSR microcomb depends on the mode-crossing location, this directly modifies the conditions necessary for the first optical parametric oscillator (OPO) threshold, subsequently affecting the microcomb’s repetition rate continuously.
To illustrate the transition between oscillating and non-oscillating regimes of the below-threshold light, we continuously tuned the resonator. By varying the voltage applied to the heaters, we observed that repetition rate and the supermode oscillations change, reaching to a non-oscillatory regime as illustrated in Fig.4(a) and Fig.4(b). Figure 4(c) presents the RF spectrum of the oscillating supermode at threshold exhibiting a reduction in the oscillation frequency to below 20 MHz, in comparison with the previous range of 100-150 MHz shown in Fig. 3(c). Further increasing the voltage leads to the complete elimination of the oscillation rate to zero within a finite temperature range, all without the need for fine-tuning (Fig. 4(d)). Increasing the voltage beyond this point resumes oscillations at approximately 20 MHz, as depicted in Fig. 4(e), showcasing the different oscillation regimes achieved through tuning.
The three oscillation regimes—fast, slow, and non-oscillating—are predicted through numerical investigations of the lattice dynamics spectrum of supermodes (Methods 4.6). Notably, the oscillation’s cessation over a finite temperature range is unexpected, suggesting that even when the APT-symmetry is not fully satisfied, but close to being satisfied, supermodes with real eigenvalues (non-oscillating) remain present. This behavior agrees with classical PT-symmetric system, that exhibit completely real eigenvalues even in non-ideal experimental conditions [37]. Our numerical investigations, detailed in Methods 4.6, indicate that this robustness is supported by our theoretical model and reveal that non-oscillatory states only begin to oscillate after a significant deviation from APT-symmetric conditions. This finding demonstrates that non-oscillatory multimode squeezed states produced by dissipative Kerr combs can be reliably maintained without stringent control over system parameters.
Figure 4: Tuning to zero oscillations of the supermodes of vacuum fluctuations a.-b. Adjusting the voltage of the micro-heater modifies the Kerr microcomb and transitions the supermode’s oscillations from non-zero (a) to zero (b). c.-d. Experimental tuning of supermode oscillations from non-zero (c) to zero (d) using micro-heaters. The RF spectrum plots are presented above the threshold, showing that the parametric gain peaks precisely at zero, as indicated by the single peak. e. The supermode oscillations as a function of micro-heater voltage demonstrate a discontinuous change from rapid oscillations to slow oscillations, ultimately leading to zero-oscillations. f. RF spectrum below the threshold in a transition region between oscillating and non-oscillating supermodes, illustrating the coexistence of both states with comparable parametric gain.
The behavior of non-oscillatory supermodes can be further understood experimentally by observing the transition between a dominant oscillating supermode to a non-oscillating supermode. The transition is not continuous in frequency but includes the two supermodes having similar parametric gain. Figure 4(f) shows exactly that - the below threshold anti-squeezing of two co-existing supermodes at the transition point. At this transition point, small changes to the temperature or input power result in either the oscillating supermode going above threshold, or the non-oscillating going above threshold. If the non-oscillating supermode reaches threshold, additional beat notes would not form as in 3(b) but a single Kerr comb with repetition rate of 1-FSR (See Section 3 in the Supplementary Material for additional experimental measurements of this phenemenon).
@@ 3 Discussion
In this work, we experimentally and theoretically investigated the lattice dynamics of optical multimode vacuum fluctuations generated by spontaneously occurring dissipative Kerr combs. We successfully resolved the RF spectrum of an oscillating supermode that spans 12 resonances. We further explored these oscillations to demonstrate their impact on the secondary threshold process, specifically the generation of classical beat notes after the threshold is exceeded. Notably, we were able to tune the supermode using micro-heaters, effectively extinguishing its oscillation completely within a finite temperature range. We elucidate this phenomenon from the perspective of non-Hermitian lattice dynamics and lattice symmetries.
Examining these fundamental properties, which connect dispersion engineering, emerging phenomena, and multimode vacuum fluctuations, paves the way for observing and engineering all-optical bosonic networks which exhibit multimode quantum squeezing and multimode quantum entanglement. Such phenomena on chip can potentially be generated by injecting or probing various pulses and multimode classical optical drives. Addressing and controlling the dynamics of supermodes is important for utilizing multimode squeezing of Kerr combs in information processing [50, 51] and sensing [52].
Oscillations and other forms of dynamical behavior can naturally arise in applications involving multimode squeezing in cavities, presenting challenges for harnessing squeezing in quantum computing or interferometry below the shot noise limit [15]. As mentioned in section 4 of the Supplementary information , a double-peak structure can directly reduce the level of squeezing measured with a multi-tone local oscillator. Thus, controlling these dynamics can help mitigate unwanted parasitic effects. Conversely, shaping the RF spectrum has proven a valuable tool for leveraging squeezing in interferometry [53]. Consequently, the techniques developed in this work for controlling this spectrum may prove advantageous for similar applications in multimode on-chip setups.
Moreover, Kerr nonlinearity allows to realize additional exotic states characterized by nontrivial artificial gauge fields, such as topological phenomena [54, 55]. Other pumping schemes may lead to non-trivial gauge fields in the quadrature space, resulting in quadrature dimers and non-reciprocal quadrature flow [7, 1, 8]. Our examination of comb formation and the vacuum noise inherent to Kerr combs provides potential for advancements in comb technology, particularly for those governed by quantum noise, which sets the ultimate limits on comb stability. The concepts articulated here form a bridge that enables exploitation of the rich non-Hermitian lattice physics and the nonlinear effects of Kerr combs to engineer complex quantum resources on chip. The lattice dynamics generated by the Kerr effect, as demonstrated in this work, are also relevant for ultrashort pulses in micro-resonators which also exhibit complex multimode bosonic network dynamics. These phenomena could manifest as below-threshold entangled supermodes, offering new opportunities for controlling quantum light.